The SAT scores for students are normally distributed with a mean of 1100 and a standard deviation of 210. What is the probability that a sample of 90 students will have an average score between 1050 and 1120? Round your answer to 3 decimal places.
Answer #1

Answer:

Given,

P(1050 < X < 1120) = P((1050 - 1100)/(210/sqrt(90)) < (x-u)/(s/sqrt(n)) < (1120 - 1100)/(210/sqrt(90)))

= P(-2.26 < z < 0.90)

= P(z < 0.90) - P(z < - 2.26)

= 0.8159398 - 0.0119106 [since from z table]

= 0.804